OpenStudy (arianna1453):

I Will Fan and Give Medals! Calculus Help!!! At what value(s) of x does f(x) = x^4 – 8x^2 have a critical point where the graph changes from decreasing to increasing?

OpenStudy (arianna1453):

@TrojanPoem

OpenStudy (trojanpoem):

Find the derivative of this function f'(x) = 4x^3 - 16x Let f'(x) = 0 4x(x^2 - 4) = 0 x^2 - 4 =0 , x = 0 x = + or - 2 or x = 0 Decreasing ]-inf , -2[ + ]0, 2[ Increasing ]-2, 0 [ + ]2, inf[

OpenStudy (trojanpoem):

Critical points : +2 , -2 , 0

OpenStudy (arianna1453):

So 0 and 2 only would be the answer?

OpenStudy (trojanpoem):

yEAH